Prove the following trigonometric identity
(tan x - sin x) / sin^3 x = 1 / cos x (1 + cos x)

Answer by filboi(35)

You can start working on the left side to get the right side of the equation.
(tan x - sin x)/(sin^3 x)
=(sin x/cos x - sin x)/sin^3 x
=(sin x -sin x cos x)/cos x * sin^3 x
Factor the numerator by sin x and split the denominator's sin^3 into sin^2 * sin x:
=sin x(1-cos x)/(cos x * sin x * sin^2 x)
=(1 - cos x)/cos x(1-cos^2 x)
Factor the denominator:
=(1 - cos x)/cos x * (1 + cos x)(1 - cos x)
=1/cos x(1 +cos x)